# SS4Rec: Continuous-Time Sequential Recommendation with State Space Models

### v1

@@ 4.5. Ablation Study
We conduct a series of ablation studies to examine each component’s contributions to the proposed architecture. Table 5 shows the performance of our default method and its five variants. The variants and their respective effects are detailed as follows:
  • Ignore time input Without the dependence on time interval sequence $\mathbf{\Delta T}$, the time-aware SSM layer depends only on item embeddings, i.e., $\mathbf{\Delta}=\mathbf{s}\in\mathbb{R}^{P}$. It reverts to the same neglect of temporal interval modeling as SASRec and others. A notable decline in the performance of SS4Rec illustrates that time input has a great effect on recommendation. Yet, its performance still surpasses most of the baselines in Table 3 and is comparable to Mamba4Rec, demonstrating the inherent advantages of the SSM architecture.
@@
4.6. Efficiency Analysis
Many previous studies have compared the time complexity of various architectures. In the ideal case of training, SSMs are generally considered to have linear time complexity, while attention-based models are often considered to have squared complexity, as shown in Table 6.
Table 6. Time Complexity. $L$ and $d$ represent the sequence length and the dimension of hidden layer respectively.
@@ 4.6. Efficiency Analysis
In practical experiments, we found that the Time and Space cost of SS4Rec is larger than that of SASRec. In time efficiency, the different discretization steps in the training phase lead to an augmented computational burden. In terms of spatial complexity, the storage of the timestamps, time intervals, and coefficient matrices after discretization with different step sizes requires much memory. A beneficial aspect is that all these costs are linearly related to the sequence length and not squared. Therefore, to compare the methods’ efficiency, we conduct experiments on longer sequence lengths, and Figure 4 reflects the variation of the ratio of SS4Rec/SASRec cost. As the length of the sequence increases, the difference in computational complexities between linear and squared becomes more apparent. Both temporal and spatial resource consumption of SASRec grows at a faster rate compared to SS4Rec. This implies that the implementation of SS4Rec in recommendation systems not only leads to enhanced performance but also exhibits greater efficiency.
[…]
Figure 4. Comparison of Practical Time/Space Cost.
4.7. Performance on Different Lengths of Sequences
Figure 5 presents the performance of SS4Rec on sequences with different lengths, compared with two representative recommendation models, attention-based model SASRec and SSM-based model Mamba4Rec. The sequences in KuaiRec(Big) dataset are divided into four groups according to their length, including (0,200], (200,400], (400,600], and (600,800], with the amount of data in each group decreasing in order.
@@ 4.7. Performance on Different Lengths of Sequences
[…]
Figure 5. Experiments of Different Lengths on KuaiRec(Big) dataset.
5. Conclusion and Future Work
@@
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China (Grants No. 62171391). Shaorong Fang and Tianfu Wu from Information and Network Center of Xiamen University are acknowledged for the help with the GPU computing.

### v2

@@ IV-E Ablation Study
The variants and their respective effects are detailed as follows:
  • Ignore time input Without the dependence on time interval sequence $\mathbf{\Delta T}$, the time-aware SSM layer depends only on item embeddings, i.e., $\mathbf{\Delta}=\mathbf{s}\in\mathbb{R}^{P}$. It reverts to the same neglect of temporal interval modeling as SASRec and others. A notable decline in the performance of SS4Rec illustrates that time input has a great effect on recommendation. Yet, its performance still surpasses most of the baselines in Table III and is comparable to Mamba4Rec, demonstrating the inherent advantages of the SSM architecture.
@@
IV-F Efficiency Analysis
Many previous studies have compared the time complexity of various architectures. In the ideal case of training, SSMs are generally considered to have linear time complexity, while attention-based models are often considered to have squared complexity, as shown in Table VI.
TABLE VI: Time Complexity. $L$ and $d$ represent the sequence length and the dimension of hidden layer respectively.
@@ IV-F Efficiency Analysis
In practical experiments, we found that the Time and Space cost of SS4Rec is larger than that of SASRec. In time efficiency, the different discretization steps in the training phase lead to an augmented computational burden. In terms of spatial complexity, the storage of the timestamps, time intervals, and coefficient matrices after discretization with different step sizes requires much memory. A beneficial aspect is that all these costs are linearly related to the sequence length and not squared. Therefore, to compare the methods’ efficiency, we conduct experiments on longer sequence lengths, and Fig. 4 reflects the variation of the ratio of SS4Rec/SASRec cost. As the length of the sequence increases, the difference in computational complexities between linear and squared becomes more apparent. Both temporal and spatial resource consumption of SASRec grows at a faster rate compared to SS4Rec. This implies that the implementation of SS4Rec in recommender systems not only leads to enhanced performance but also exhibits greater efficiency.
Figure 4: Comparison of practical time/space cost.
IV-G Performance on Different Lengths of Sequences
Fig. 5 presents the performance of SS4Rec on sequences with different lengths, compared with two representative recommendation models, attention-based model SASRec and SSM-based model Mamba4Rec. The sequences in KuaiRec(Big) dataset are divided into four groups according to their length, including (0,200], (200,400], (400,600], and (600,800], with the amount of data in each group decreasing in order.
@@ IV-G Performance on Different Lengths of Sequences
Figure 5: Experiments of different lengths on KuaiRec(Large) dataset.
V Conclusion and Future Work